OpenStudy (anonymous):

Could you check my answer? @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how are you getting 502.8 ?

jimthompson5910 (jim_thompson5910):

I see now, thanks

jimthompson5910 (jim_thompson5910):

What I would do is replace each '5' with '9*5' and replace every '11' with '9*11' we do this because each dimension is 9 times longer then recompute the surface area of the new figure

jimthompson5910 (jim_thompson5910):

keep in mind that 5^2 will turn into (9*5)^2

OpenStudy (anonymous):

\[2 \times 22 \div 7 \times 9*5 + 2 \times 22 \div 7 \times 5 \times 9 * 11 = ?\]

jimthompson5910 (jim_thompson5910):

you forgot to square the "9*5" portion

OpenStudy (anonymous):

Oh that's was meant to be (9*5)^2

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

also, you forgot to replace the second '5' with '9*5'

OpenStudy (anonymous):

\[2 \times 22 \div 7 \times (9*5)^2 + 2 \times 22 \div 7 \times 9*5 \times 9 * 11 = ? \]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now compute that

OpenStudy (anonymous):

I got a SUPER HUGE # !

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

178218183.7

OpenStudy (anonymous):

>.<

jimthompson5910 (jim_thompson5910):

hmm, strange, it's not supposed to be that big

jimthompson5910 (jim_thompson5910):

well anyways, I did 9*5 = 45 9*11 = 99 so that means... 2*(22/7)*45^2+2*(22/7)*45*99 = 40,731.42857 (approximately)

jimthompson5910 (jim_thompson5910):

yeah we're rounding to the nearest whole number

jimthompson5910 (jim_thompson5910):

check to make sure you get the same (and try to figure out what went wrong)

OpenStudy (anonymous):

Yes, now I got that answer!

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome